CS 791 - Network Optimization

K. Subramani LDCSEE, West Virginia University, Morgantown, WV ksmani@csee.wvu.edu

1 Policies and Procedures

(i) All homework assignments must be turned in on the prescribed date. (ii) Homework assignments must be written up in LaTeX. (iii) A maximum of 5 bonus points is reserved for class performance. (iv) Each presentation should be self-contained and include the following information:

(a) Statement of Problem,

(b) Related Work,

(c) Current contribution,

(d) Open problems.

(v) If your presentation can be converted into publishable work, you will be awarded an A, regardless of your homework performance.

(vi) The standard Network Optimization conferences are: (a) ACM Symposium on Discrete Algorithms.

(b) Integer Programming and Combinatorial Optimization. (vii) The standard Network Optimization journals are:

(a) Networks.

(b) Journal of Graph Theory and Applications.